OpenStudy (anonymous):

how can i find the volume of a tetrahedron given the altitude of 10cm

OpenStudy (rose21):

altitude = root(2/3) * a 10 = root(2/3) * a 100=2/3 a^2 a= 5 root 6 V = root(2)/12 * a^3 V = 216.5

OpenStudy (rose21):

does this help u ?

OpenStudy (anonymous):

can you explain it please its my exam tom

OpenStudy (anonymous):

what is a in that thing?

OpenStudy (rose21):

i am not really sure how to explain it more i am only 14

OpenStudy (rose21):

i am really sorry

OpenStudy (anonymous):

what is 12 on the volume? what is the actual formula. please reply

OpenStudy (anonymous):

what's the formula